# Tree Embedding in High Dimensions:

### v1

@@ 4 Dynamic and MPC Implementations of Algorithm˜1
The notion of consistent hashing was introduced by [CJK+22] which is deterministic, and the notion in Definition˜4.1 is a relaxed randomized version which was introduced by [FJL+25]. It turns out that the expected guarantee of consistency already suffices for our applications. For example, if we have a $(\Gamma,\Lambda)$-hash, then $\mathbb{E}[|\tilde{B}^{P}_{i}(p,r_{i})|]\leq\Lambda$ (as in Line 1). We state our theorems for dynamic and MPC tree embedding with respect to a general $(\Gamma,\Lambda)$-hash, as follows. Their proof can be found in Sections˜4.2 and 4.1.
@@
We discuss the concrete hashing bounds that can be plugged into Theorems˜4.2 and 4.3. [FJL+25] gave a hash function that satisfies the requirements of theorem˜4.2, for $\Lambda=2^{O(d/\Gamma^{2/3})}\cdot\operatorname{poly}(d)$. Moreover, in [CJK+22] a hashing that satisfies the requirements of Theorem˜4.3 was presented and $\Lambda=2^{O(d/\Gamma)}\cdot\operatorname{poly}(d)$. These two results are (re)stated as follows.
@@
The second property in Lemma˜4.4 is not explicitly stated, but can be concluded from [FJL+25]. We do not claim novelty for it and elaborate on its execution in Appendix˜A.
@@
Although these bounds have an exponential in $d$ dependence in $\Lambda$, they are actually enough for implying the main theorems we state in Section˜1.1, since one can without loss of generality applying a standard Johnson-Lindenstrauss transform [JL84] to reduce $d$ to $d=O(\log n)$, and this only increases the distortion bound by a constant factor, which we can afford. This way, a bound of $2^{d/t}$ would lead to $n^{1/t}$ (for $t\geq 1$). More specifically, such a transform $\phi:\mathbb{R}^{d}\to\mathbb{R}^{m}$ for $m=O(\log n)$ is data-oblivious, and can compute for each input point $x\in\mathbb{R}^{d}$ its image $\phi(x)$ in $O(dm)$ time and space, which can be trivially implemented in dynamic or MPC settings.
4.1 Proof of Theorem˜4.2: Dynamic Implementation
@@
Our dynamic algorithm will maintain the tree embedding which is the output of the static Algorithm˜1. Hence, the main idea of the dynamic algorithm is to correctly maintain the labels $\ell_{p}^{(i)}$ for all points $p\in P$, for all levels $i\in[m]$. By Lemma˜3.1 the tree embedding defined by the labels will then have distortion $O(\Gamma\cdot\log\Gamma\cdot\log n)$, implying the correctness of the dynamic algorithm.
@@
Item˜3 essentially says that on all levels $i\in[m]$ for any hash value $x$ with at least one point assigned to it from $P$ by $\varphi_{i}$ we maintain all points in $P$ with $\varphi_{i}$ value $x$. We note that we use standard dynamic set representations to maintain $\varphi_{i}^{-1}(x)\cap P$, and use min-heaps to maintain values $\pi_{\min}(\varphi_{i}^{-1}(x)\cap P)$.
@@
Now we turn to the more challenging task of point insertions. Under point insertions, our dynamic algorithm will exactly maintain the output of the static Algorithm˜1. As mentioned, this only requires us to correctly maintain the labels $\ell_{p}^{(i)}$ for all $p\in P,i\in[m]$. The details of the algorithm are presented in Algorithm˜2.
@@ 17            Update $\ell_{p^{\prime}}^{(i)}\leftarrow\pi(p)$
At insertion of point $p\in\mathbb{R}^{d}$, we first draw $\pi(p)$ and calculate $\varphi_{i}(p)$ for $i\in[m]$, i.e. the auxiliary variables in Item˜1 and Item˜2.
At each scale $i$, the algorithm Algorithm˜2 then proceeds in two steps. First, we find the labels of the inserted point $p$ (lines 4-5), and then we update the labels of other points if necessary (lines 6-9).
@@
Running Time Analysis.
Now we discuss the running time guarantees of the algorithm.
@@
Algorithm˜2 runs in $\tilde{O}(\Lambda+d)$ amortized update time.
@@
Note that visiting one specific point takes $O(1)$ time. Consider when point $p\in P$ might be visited. This occurs when for some hash value $x\in\varphi_{i}(P\cap B(p,r_{i}))$ the value $\pi_{\min}(\varphi^{-1}_{i}(x)\cap P)$ is populated (in case it is the first point insertion associated with that hash value) or updated. By the definition of the hash function, we know that $\mathbb{E}[|x\in\varphi_{i}(P\cap B(p,r_{i}))|]\leq\Lambda$ and by Lemma˜4.7, each such hash value may change its minimal element according to $\pi$ in $P$ at most $O(\log n)$ times in expectation over $n$ point insertions. Hence, each point is visited at most $\tilde{O}(\Lambda)$ times in expectation over the series of $n$ updates. Combining all running-time components, we achieve the desired $\tilde{O}(\Lambda+d)$ amortized update time for insertion operations, with the additional $\tilde{O}(d)$ term resulting from the Johnson-Lindenstrauss transform.
The case of point deletions is now trivial. First, the deactivation of a node takes $O(1)$ time. Since we re-computing the embedding every $\Theta(n)$ updates by inserting all existing points into an empty input, we ensure that at all times the input only contains $O(n)$ inactive points. Since the total update time of the algorithm over a sequence of $n$ points insertions is $\tilde{O}((d+\Lambda)\cdot n)$, the cost of these recomputation amortized over the $\Theta(n)$ updates between rebuilds works out to be an additive $\tilde{O}(d+\Lambda)$ factor, as desired. ∎
@@
As discussed, our algorithm implicitly maintains the tree embedding by storing labels $\ell_{p}^{(i)}$ for all points $p\in P$ at all levels $i\in[m]$. Hence, the tree embedding can be queried by simply reporting these labels for each point. However, it is useful to examine more carefully the structural changes to the tree embedding that arise from label updates. This is especially relevant to applications of our dynamic tree embedding, as presented in Section˜5.
@@
When computing the labels of $p$ (lines 4–5 of Algorithm˜2), notice that at each level $i$, the inserted point either receives a new label that no other point at level $i$ has, or a label that some other points already have. Hence, exactly one new leaf and its connecting path may appear during each insertion; that is, an update of Type 2) occurs in this part.
On the other hand, changing the labels of existing points at level $i$ (lines 6–9 of Algorithm˜2) corresponds to certain subsets of points on some of the layers $i\in[m]$ changing their labels to correspond to the newly inserted point. We We will now show that this kind of update, where nodes switch parents within the tree, can be simulated by Type 1) and Type 2) updates.
@@
It remains to argue that the simulation of edge changes through point deactivations and insertions does not lead to a significant blowup in update time or tree size. Consider the label of any point on any level $i\in[m]$. It will always correspond to some $\pi_{\min}$ value for some incremental set of at most $n$ uniform $[0,1]$ random variables. Hence, by Lemma˜4.7 it will be updated $O(\log n)$ times in expectation over the series of $n$ updates.
@@
4.2 Proof of Theorem˜4.3: MPC Implementation
@@
Proof of Theorem˜4.3.
We only need to show how to simulate Algorithm˜1 in MPC, and the distortion bound follows immediately from Lemma˜3.1.
For steps other than Line 1, the value $\beta$ may be generated on a leader machine and then use a standard broadcast (see e.g., [Gha19]) to make it available to every machine, the $\pi$ value may be generated independently on every machine. The for-loop may be run in parallel, and $r_{i}$ may be computed locally. The hash is deterministic and data-oblivious, so every machine can have the same hash without communication, and within the space budget by Lemma˜4.5.
@@ 5 Applications: New Dynamic Algorithms in High Dimension
Combining these results with the dynamic algorithm for tree embedding (Theorem˜4.2), we obtain dynamic algorithms for $k$-median (Corollary˜5.1), bipartite matching (Corollary˜5.5), geometric transport (corollary˜5.11) and general matching (corollary˜5.14) w.r.t. Euclidean metric.
Throughout this section, we will be referring to the embedding tree as $T$ with $m\in\tilde{O}(1)$ layers. To stay consistent with the notations of section˜3, the root will be on level $1$ and leaves will be on level $m$. For sake of simplicity, we will assume that the distance of a leaf and its parent is $1$, and hence the distance of a node on level $l$ and $l+1$ is $2^{m-l}$. We will refer to the level of point $p$ in the tree as $\ell(v)\in[m]$. We will refer to the set of input points represented by the leaves in the sub-tree of node $v$ as $P(v)$. For any node $v\in T$, we will refer to the set of nodes in the sub-tree of $v$ as $T_{v}$.
@@ 5.1 $k$-Median
To prove the above result, we first present a dynamic algorithm that maintains an exact $k$-median w.r.t. tree metric $T$ under operations described in Section˜5. corollary˜5.1 is a direct implication of lemma˜5.2, theorem˜4.2 and lemma˜4.4.
@@
In our dynamic implementation, each node $v$ of $T$ at level $\ell(v)$ maintains a max heap $H(v)$ with elements of form $(\operatorname{center}(u),\operatorname{value}(u))$ for all $u\in C(v)$, and $\operatorname{value}(u)$ is defined as $\text{benf}(u)-P_{\ell(v)+1}(u)\cdot 2^{m-\ell(v)}$, sorted based on the second value. Hence, if $(u,x)$ is the first element of $H(v)$ then by ˜5.3 $\text{benf}(v)=x+\text{pes}(v)$ and $\operatorname{center}(v)=u$.
Leaf deletions.
We first describe how to maintain the max-heaps $H(v)$ and $\text{benf}(v),\operatorname{center}(v)$ values when leaf $v\in T$ is deactivated, that is under Type 2) updates.
@@ 8 $\operatorname{center}(u)\leftarrow w$
Path insertions.
Now we describe the operations required to maintain the same values under Type 1) updates, that is when a path $\Pi$ is inserted connecting leaf $v$ to some already existing node $r_{v}$. Note that the process is split to two parts only because for the new set of nodes and $r_{v}$ a new element has to be added to their max-heaps.
@@ 7 $\text{benf}(u)\leftarrow\text{benf}(u^{\prime})-2^{m-\ell(u)}+\text{pes}(u)$
8 end for
9Let $\Pi^{\prime}$ be a path from $r_{v}$ to root $r$
@@ 14 $\operatorname{center}(u)\leftarrow w$
First, observe that when leaf $v$ is either inserted or deleted, $\operatorname{center}$ and benf values may only change for nodes of $T$ whose sub-tree contains $v$, that is along a path $\Pi$ from $v$ to the root. For leaf $v$, the algorithm updates these values trivially. For any non-leaf $u\in\Pi$, the algorithm in a bottom up manner first restores the correctness of heap $H(u)$ (whose only value which has to be updated is the one corresponding to the single node in $C(u)\cap\Pi$). Afterwards, it updates $\text{benf}(u)$ and $\operatorname{center}(u)$ correctly due to ˜5.3.
@@
Queries.
In order to answer queries, its sufficient to maintain for all $v\in T$ a list of all leafs in $T_{v}$ which are part of the $k$-median solution. This information can be updated in $\tilde{O}(1)$ time per update to the $k$-median solution through walking along a leaf-root path.
@@ 5.2 Bipartite Euclidean Matching
In particular, our algorithm achieves $O_{\epsilon}(\log n)$ approximation in $\tilde{O}(n^{\epsilon}+d)$ update time for any constant $\epsilon>1$ or $\tilde{O}(1)$-approximation in $\tilde{O}(d)$ update time. Similarly to the $k$-median problem, using the dynamic tree embedding algorithm of Theorem˜4.2, in order to obtain a dynamic bipartite Euclidean matching algorithm, it is sufficient to solve the problem on dynamic 2-HST-s under the specific updates defined by the theorem (Type 1) and 2)). corollary˜5.5 is a direct implication of theorem˜4.2, lemma˜4.4 and lemma˜5.6.
@@
Static Algorithm.
We will first describe a simple static algorithm computing such a matching in time $\tilde{O}(n)$. Our dynamic implementation will exactly maintain the output of this algorithm. For all nodes $v\in T$, let $A(v)$ and $B(v)$ stand for the set of red and blue leaves in the sub-tree of $v$. Then $P(v)=A(v)\cup B(v)$ stands for the points associated with these leaves.

### v2

@@ 4 Dynamic and MPC Implementations of Algorithm 1
The notion of consistent hashing was introduced by [CJK+22] which is deterministic, and the notion in Definition 4.1 is a relaxed randomized version which was introduced by [FJL+25]. It turns out that the expected guarantee of consistency already suffices for our applications. For example, if we have a $(\Gamma,\Lambda)$-hash, then $\mathbb{E}[|\tilde{B}^{P}_{i}(p,r_{i})|]\leq\Lambda$ (as in Line 1). We state our theorems for dynamic and MPC tree embedding with respect to a general $(\Gamma,\Lambda)$-hash, as follows. Their proof can be found in Sections 4.2 and 4.1.
@@
We discuss the concrete hashing bounds that can be plugged into Theorems 4.2 and 4.3. [FJL+25] gave a hash function that satisfies the requirements of Theorem 4.2, for $\Lambda=2^{O(d/\Gamma^{2/3})}\cdot\operatorname{poly}(d)$. Moreover, in [CJK+22] a hashing that satisfies the requirements of Theorem 4.3 was presented and $\Lambda=2^{O(d/\Gamma)}\cdot\operatorname{poly}(d)$. These two results are (re)stated as follows.
@@
The second property in Lemma 4.4 is not explicitly stated, but can be concluded from [FJL+25]. We do not claim novelty for it and elaborate on its execution in Appendix A.
@@
Although these bounds have an exponential in $d$ dependence in $\Lambda$, they are actually enough for implying the main theorems we state in Section 1.1, since one can without loss of generality applying a standard Johnson-Lindenstrauss transform [JL84] to reduce $d$ to $d=O(\log n)$, and this only increases the distortion bound by a constant factor, which we can afford. This way, a bound of $2^{d/t}$ would lead to $n^{1/t}$ (for $t\geq 1$). More specifically, such a transform $\phi:\mathbb{R}^{d}\to\mathbb{R}^{m}$ for $m=O(\log n)$ is data-oblivious, and can compute for each input point $x\in\mathbb{R}^{d}$ its image $\phi(x)$ in $O(dm)$ time and space, which can be trivially implemented in dynamic or MPC settings.
4.1 Proof of Theorem 4.2: Dynamic Implementation
@@
Our dynamic algorithm will maintain the tree embedding which is the output of the static Algorithm 1. Hence, the main idea of the dynamic algorithm is to correctly maintain the labels $\ell_{p}^{(i)}$ for all points $p\in P$, for all levels $i\in[m]$. By Lemma 3.1 the tree embedding defined by the labels will then have distortion $O(\Gamma\cdot\log\Gamma\cdot\log n)$, implying the correctness of the dynamic algorithm.
@@
Item 3 essentially says that on all levels $i\in[m]$ for any hash value $x$ with at least one point assigned to it from $P$ by $\varphi_{i}$ we maintain all points in $P$ with $\varphi_{i}$ value $x$. We note that we use standard dynamic set representations to maintain $\varphi_{i}^{-1}(x)\cap P$, and use min-heaps to maintain values $\pi_{\min}(\varphi_{i}^{-1}(x)\cap P)$.
@@
Now we turn to the more challenging task of point insertions. Under point insertions, our dynamic algorithm will exactly maintain the output of the static Algorithm 1. As mentioned, this only requires us to correctly maintain the labels $\ell_{p}^{(i)}$ for all $p\in P,i\in[m]$. The details of the algorithm are presented in Algorithm 2.
@@ 17            Update $\ell_{p^{\prime}}^{(i)}\leftarrow\pi(p)$
At insertion of point $p\in\mathbb{R}^{d}$, we first draw $\pi(p)$ and calculate $\varphi_{i}(p)$ for $i\in[m]$, i.e. the auxiliary variables in Item 1 and Item 2.
At each scale $i$, the algorithm Algorithm 2 then proceeds in two steps. First, we find the labels of the inserted point $p$ (lines 4-5), and then we update the labels of other points if necessary (lines 6-9).
@@
Run Time Analysis.
Now we discuss the run-time guarantees of the algorithm.
@@
Algorithm 2 runs in $\tilde{O}(\Lambda+d)$ amortized update time.
@@
Note that visiting one specific point takes $O(1)$ time. Consider when point $p\in P$ might be visited. This occurs when for some hash value $x\in\varphi_{i}(P\cap B(p,r_{i}))$ the value $\pi_{\min}(\varphi^{-1}_{i}(x)\cap P)$ is populated (in case it is the first point insertion associated with that hash value) or updated. By the definition of the hash function, we know that $\mathbb{E}[|x\in\varphi_{i}(P\cap B(p,r_{i}))|]\leq\Lambda$ and by Lemma 4.7, each such hash value may change its minimal element according to $\pi$ in $P$ at most $O(\log n)$ times in expectation over $n$ point insertions. Hence, each point is visited at most $\tilde{O}(\Lambda)$ times in expectation over the series of $n$ updates. Combining all running-time components, we achieve the desired $\tilde{O}(\Lambda+d)$ amortized update time for insertion operations, with the additional $\tilde{O}(d)$ term resulting from the Johnson-Lindenstrauss transform.
The case of point deletions is now trivial. First, the deactivation of a node takes $O(1)$ time. Since we re-computing the embedding every $\Theta(n)$ updates by inserting all existing points into an empty input, we ensure that at all times the input only contains $O(n)$ inactive points. Since the total update time of the algorithm over a sequence of $n$ points insertions is $\tilde{O}((d+\Lambda)\cdot n)$, the cost of these recomputation amortized over the $\Theta(n)$ updates between rebuilds works out to be an additive $\tilde{O}(d+\Lambda)$ factor, as desired.
@@
As discussed, our algorithm implicitly maintains the tree embedding by storing labels $\ell_{p}^{(i)}$ for all points $p\in P$ at all levels $i\in[m]$. Hence, the tree embedding can be queried by simply reporting these labels for each point. However, it is useful to examine more carefully the structural changes to the tree embedding that arise from label updates. This is especially relevant to applications of our dynamic tree embedding, as presented in Section 5.
@@
When computing the labels of $p$ (lines 4–5 of Algorithm 2), notice that at each level $i$, the inserted point either receives a new label that no other point at level $i$ has, or a label that some other points already have. Hence, exactly one new leaf and its connecting path may appear during each insertion; that is, an update of Type 2) occurs in this part.
On the other hand, changing the labels of existing points at level $i$ (lines 6–9 of Algorithm 2) corresponds to certain subsets of points on some of the layers $i\in[m]$ changing their labels to correspond to the newly inserted point. We We will now show that this kind of update, where nodes switch parents within the tree, can be simulated by Type 1) and Type 2) updates.
@@
It remains to argue that the simulation of edge changes through point deactivations and insertions does not lead to a significant blowup in update time or tree size. Consider the label of any point on any level $i\in[m]$. It will always correspond to some $\pi_{\min}$ value for some incremental set of at most $n$ uniform $[0,1]$ random variables. Hence, by Lemma 4.7 it will be updated $O(\log n)$ times in expectation over the series of $n$ updates.
@@
4.2 Proof of Theorem 4.3: MPC Implementation
@@
Proof of Theorem 4.3.
We only need to show how to simulate Algorithm 1 in MPC, and the distortion bound follows immediately from Lemma 3.1.
For steps other than Line 1, the value $\beta$ may be generated on a leader machine and then use a standard broadcast (see e.g., [Gha19]) to make it available to every machine, the $\pi$ value may be generated independently on every machine. The for-loop may be run in parallel, and $r_{i}$ may be computed locally. The hash is deterministic and data-oblivious, so every machine can have the same hash without communication, and within the space budget by Lemma 4.5.
@@ 5 Applications: New Dynamic Algorithms in High Dimension
Combining these results with the dynamic algorithm for tree embedding (Theorem 4.2), we obtain dynamic algorithms for $k$-median (Corollary 5.1), bipartite matching (Corollary 5.5), geometric transport (Corollary 5.11) and general matching (Corollary 5.14) w.r.t. Euclidean metric.
Throughout this section, we will be referring to the embedding tree as $T$ with $m\in\tilde{O}(1)$ layers. To stay consistent with the notations of Section 3, the root will be on level $1$ and leaves will be on level $m$. For sake of simplicity, we will assume that the distance of a leaf and its parent is $1$, and hence the distance of a node on level $l$ and $l+1$ is $2^{m-l}$. We will refer to the level of point $p$ in the tree as $\ell(v)\in[m]$. We will refer to the set of input points represented by the leaves in the sub-tree of node $v$ as $P(v)$. For any node $v\in T$, we will refer to the set of nodes in the sub-tree of $v$ as $T_{v}$.
@@ 5.1 $k$-Median
To prove the above result, we first present a dynamic algorithm that maintains an exact $k$-median w.r.t. tree metric $T$ under operations described in Section 5. Corollary 5.1 is a direct implication of Lemma 5.2, Theorem 4.2 and Lemma 4.4.
@@
In our dynamic implementation, each node $v$ of $T$ at level $\ell(v)$ maintains a max heap $H(v)$ with elements of form $(\operatorname{center}(u),\operatorname{value}(u))$ for all $u\in C(v)$, and $\operatorname{value}(u)$ is defined as $\text{benf}(u)-P_{\ell(v)+1}(u)\cdot 2^{m-\ell(v)}$, sorted based on the second value. Hence, if $(u,x)$ is the first element of $H(v)$ then by 5.3 $\text{benf}(v)=x+\text{pes}(v)$ and $\operatorname{center}(v)=u$.
Leaf deletions. We first describe how to maintain the max-heaps $H(v)$ and $\text{benf}(v),\operatorname{center}(v)$ values when leaf $v\in T$ is deactivated, that is under Type 2) updates.
@@ 8 $\operatorname{center}(u)\leftarrow w$
Path insertions. Now we describe the operations required to maintain the same values under Type 1) updates, that is when a path $\Pi$ is inserted connecting leaf $v$ to some already existing node $r_{v}$. Note that the process is split to two parts only because for the new set of nodes and $r_{v}$ a new element has to be added to their max-heaps.
@@ 7 $\text{benf}(u)\leftarrow\text{benf}(u^{\prime})-2^{m-\ell(u)}+\text{pes}(u)$
8Let $\Pi^{\prime}$ be a path from $r_{v}$ to root $r$
@@ 13 $\operatorname{center}(u)\leftarrow w$
First, observe that when leaf $v$ is either inserted or deleted, $\operatorname{center}$ and benf values may only change for nodes of $T$ whose sub-tree contains $v$, that is along a path $\Pi$ from $v$ to the root. For leaf $v$, the algorithm updates these values trivially. For any non-leaf $u\in\Pi$, the algorithm in a bottom up manner first restores the correctness of heap $H(u)$ (whose only value which has to be updated is the one corresponding to the single node in $C(u)\cap\Pi$). Afterwards, it updates $\text{benf}(u)$ and $\operatorname{center}(u)$ correctly due to 5.3.
@@
Queries. In order to answer queries, its sufficient to maintain for all $v\in T$ a list of all leafs in $T_{v}$ which are part of the $k$-median solution. This information can be updated in $\tilde{O}(1)$ time per update to the $k$-median solution through walking along a leaf-root path.
@@ 5.2 Bipartite Euclidean Matching
In particular, our algorithm achieves $O_{\epsilon}(\log n)$ approximation in $\tilde{O}(n^{\epsilon}+d)$ update time for any constant $\epsilon>1$ or $\tilde{O}(1)$-approximation in $\tilde{O}(d)$ update time. Similarly to the $k$-median problem, using the dynamic tree embedding algorithm of Theorem 4.2, in order to obtain a dynamic bipartite Euclidean matching algorithm, it is sufficient to solve the problem on dynamic 2-HST-s under the specific updates defined by the theorem (Type 1) and 2)). Corollary 5.5 is a direct implication of Theorem 4.2, Lemma 4.4 and Lemma 5.6.
@@
Static Algorithm. We will first describe a simple static algorithm computing such a matching in time $\tilde{O}(n)$. Our dynamic implementation will exactly maintain the output of this algorithm. For all nodes $v\in T$, let $A(v)$ and $B(v)$ stand for the set of red and blue leaves in the sub-tree of $v$. Then $P(v)=A(v)\cup B(v)$ stands for the points associated with these leaves.